OpenStudy (amour.anohs):

he three finalists in the talent show are Emily, Miguel, and Valerie. Their combined score is 24. Emily and Miguel's combined score is twice that of Valerie, and Valerie scored only one more point than Miguel. How many points did Emily score? Write an equation that represents the situation?

OpenStudy (amour.anohs):

e + m + v = 24 e + m = 2v Answer :v = m + 1

candycove (candycove):

omg

OpenStudy (amour.anohs):

Lol Why You Keep Putting Omg On My Q's And R's